OpenStudy (anonymous):

A cylinder has a radius of 12 m and a height of 9 m. What is the exact volume of the cylinder?

OpenStudy (anonymous):

I think it is A

OpenStudy (anonymous):

\[volume =\pi r^2 h\]

OpenStudy (anonymous):

r=12 m h=9 m

OpenStudy (anonymous):

So if do 3.14*12*9?

OpenStudy (anonymous):

\[r^2=r*r\]

OpenStudy (anonymous):

i got 4069 o,o